### 19 Jan 2024 v1

@@ 2.1 Initial Set of Ethical Concerns
Internal and External Concerns Not only has mobile application software created ethical concerns for users in virtual spaces, but these concerns have also bled into the physical realm. Within the past few years, software companies have received criticism for influencing elections, inciting violence in the real world, and facilitating sexual assault ([4, 51, 66, 41, 54]). To identify and analyze these material matters, we consider any concerns that involve the physical world as external ethical concerns. We identified six ethical concerns as potentially external: accountability, discrimination, safety, scam, sustainability, and accessibility [^1] The scam ethical concern was added after the analysis in Section 2.4. This consideration can help developers understand the scope of the ethical concern and help them identify and mitigate it. Similarly, we consider internal ethical concerns as those that occur within the realm of the software application. For example, end users discriminate against end users within the software platform.
@@ 3.1 Types of Ethical Concerns (RQ1)
Internal and External Ethical Concerns Within the entire dataset, (1768/1919) 92% of the ethical concerns were labeled internal, meaning that there was an ethical issue within the application. Within the six categories we identified as potentially external (scam, safety, accountability, sustainability, discrimination), 80.3% (583/726) of the ethical concerns were labeled internal. Figure 4 shows the breakdown by ethical concern. The two concerns with the most external labels are scam and safety. These were mostly associated with the Uber application, with users reporting that drivers engage in sexual harassment and cause dangerous situations.
@@ 3.3 Characteristics of Ethical Concerns (RQ2)
Table 4: Examples of Internal and External Ethical Concerns in User Reviews
Ethical Concern User Review Platform
External (Safety)
Horrible customer service and they don’t follow their own community guidelines. They allow dangerous medical misinformation and animal mistreatment to flourish.
TikTok
External (Discrimination)
The right wing has started using your partner drivers to spread communal poison(islamophobia, calls for violence) and verifiably fake news( Uber delhi ). Multiple friends have had to deal with unsolicited attempts at indoctrination. You should have a public relations team look into this. Your app doesn’t have a field where I can leave descriptive negative feedback, or upload conversation recordings, and that brought me here.
Uber
Internal (Privacy)
Accesses a bunch of personal information on your phone for no good reason. Pay for the service twice. Once with your privacy and then with your money.
Uber
Internal (Accountability)
I never thought I’d see a platform so poorly managed but here we are. TikTok randomly removes videos for community guideline violations when they do not violate guidelines. These random videos are largely minorities (BIPOC, disabled, etc) commenting on or criticizing popular creators’ works. TikTok is inaccessible. They don’t have in-app captioning, making it difficult to even impossible for creators to make their content accessible. TikTok refuses to accept accountability for any of this.
TikTok
@@ 3.4 Automated Analysis of Ethical Concerns (RQ3)
Binary Classification of Internal and External Ethical Concerns For some reviews mentioning ethical concerns (3/583), it was difficult to assign an internal/external label because both labels applied or annotators were unsure about their labeling. Therefore, we left these reviews out of the data we used to train and test internal/external models. With the inclusion of only our potential external categories, detailed in 2.1, we had a dataset that contained $583583$ reviews labeled as internal and 143 reviews labeled as external. We used this data for the training and testing of the models.
@@ 5.2 Internal Validity
When labelling for addiction, we labelled reviews mentioning overuse of the app, even if no negative consequences were detailed in the review. We made this decision due to the complex nature of the ethical concern. In order for some phenomenon to be addictive, it must have a highly positive aspect to draw the user. Due to the critical nature of this ethical concern, and clear evidence of the negative health effects on users [31, 32, 57], we consider it crucial to ensure these reviews are considered in our analysis. Nevertheless, there is a potential overlabelling of addiction as a concern due to a lack of provided information from users in app reviews.

### 06 Feb 2024 v2

@@ 2.1 Initial Set of Ethical Concerns
2.1.1 Internal and External Concerns
Not only has mobile application software created ethical concerns for users in virtual spaces, but these concerns have also bled into the physical realm. Within the past few years, software companies have received criticism for influencing elections, inciting violence in the real world, and facilitating sexual assault ([4, 51, 66, 41, 54]). To identify and analyze these material matters, we consider any concerns that involve the physical world as external ethical concerns. We identified six ethical concerns as potentially external: accountability, discrimination, safety, scam, sustainability, and accessibility [^1] The scam ethical concern was added after the analysis in Section 2.4. This consideration can help developers understand the scope of the ethical concern and help them identify and mitigate it. Similarly, we consider internal ethical concerns as those that occur within the realm of the software application. For example, end users discriminate against end users within the software platform.
@@ 3.1 Types of Ethical Concerns (RQ1)
Internal and External Ethical Concerns Within the entire dataset, (1768/1919) 92% of the ethical concerns were labeled internal, meaning that there was an ethical issue within the application. Within the six categories we identified as potentially external (scam, safety, accountability, sustainability, discrimination), 80.3% (583/726) of the ethical concerns were labeled internal. Figure 4 shows the breakdown by ethical concern. The two concerns with the most external labels are scam and safety. These were mostly associated with the Uber application, with users reporting that drivers engage in sexual harassment and cause dangerous situations. See Table 2 for examples.
Table 2: Examples of Internal and External User Reviews
Ethical Concern External Internal
Safety
Awful customer service and they don’t follow their own community guidelines. They allow dangerous medical misinformation and animal mistreatment to flourish.
I keep getting a message that the Alexa app has crashed without using the app. This is actually dangerous when I’m using a navigation app while driving and the Alexa message shows up and blocks my view of the screen so I can’t see the map until I dismiss the Alexa message.
Scam
After I started posting videos on TikTok, I began to receive scam phone calls.
It’s difficult to use this app even though it should be easy. When I booked the ride today, the fare estimate was 39Rs but I was charged 100Rs.
Accountability
Awful service. My Uber driver refused to come and did not answer my call. He didn’t even cancel the ride. It was late at night and I was standing on road where there were no other options. Pathetic customer service…
I really like the app but unfortunately many creators aren’t held accountable for violating community guidelines even after users report them.
Sustainability
Uber should be made to pay an environmental emissions tax in every country it operates in without increasing prices or decreasing payouts to drivers.
Since new update, the app is draining my phone battery where I have to charge the phone multiple times every day. I disabled the app and my battery is back to lasting all day. Please fix. Thank you.
Discrimination
The right wing has started using your partner drivers to spread communal poison (islamophobia, calls for violence) and verifiably fake news. Your app doesn’t have a field where I can leave descriptive negative feedback, or upload conversation recordings, and that brought me here.
I have never received a single job contact from this site, nor from this company that I have often applied too, I am sure like all the tech companies you do not hire African Americans, yet I still use the app because it’s ”professional”.
@@ 3.4 Automated Analysis of Ethical Concerns (RQ3)
Binary Classification of Internal and External Ethical Concerns For some reviews mentioning ethical concerns (3/583), it was difficult to assign an internal/external label because both labels applied or annotators were unsure about their labeling. Therefore, we left these reviews out of the data we used to train and test internal/external models. With the inclusion of only our potential external categories, detailed in 2.1.1, we had a dataset that contained $583583$ reviews labeled as internal and 143 reviews labeled as external. We used this data for the training and testing of the models.
@@ 4.4 Implications
The internal and external classification provides the developer with more information regarding the extent of the issue. Once an ethical concern becomes external, it may become a higher priority, as it may present a physical threat to the user(s). In these situations, there is potential risk for serious harm for the user, litigation, and damage to the platform’s reputation. As such, it may be necessary to involve other departments (legal, public relations) while crafting a solution. It also may be critical for software professionals to contact proper authorities or otherwise quickly intervene to ensure the safety of their users (consider situations regarding sexual assault, self-harm, or other acts of physical violence). It could also therefore provide practitioners with information regarding the scheduling of various tasks.
@@ 5.2 Internal Validity
When labelling for addiction, we labelled reviews mentioning overuse of the app, even if no negative consequences were detailed in the review. We made this decision due to the complex nature of the ethical concern. In order for some phenomenon to be addictive, it must have a highly positive aspect to draw the user. Due to the critical nature of this ethical concern, and clear evidence of the negative health effects on users [31, 32, 57], we consider it crucial to ensure these reviews are considered in our analysis. Nevertheless, there is a potential overlabelling of addiction as a concern due to a lack of provided information from users in app review due to its vernacular use as hyperbole (‘I’m addicted’ sometimes just means ‘I love it’). This may not just inflate frequency but the user review score.